my problem says 'Sarah bought a laptop for $1800. The total cost, including tax, came to $1890. What is the tax rate?' I know there is a certain formula to solve it, I'm just not sure what it is.
Found 2 solutions by nerdybill, robertb:
Answer by nerdybill(7384)

Sarah bought a laptop for $1800. The total cost, including tax, came to $1890. What is the tax rate?
.
Let x = tax rate
then
1800 + 1800x = 1890
1800(1 + x) = 1890
1 + x = 1890/1800
x = 1890/1800 - 1
x = 1.05 - 1
x = .05
or
x = 5%


Answer by robertb(5830)